Twenty neurons synapse with a single receptor neuron. Fifteen of the twenty neurons release neurotransmitters that produce EPSPs in the post-synaptic cell, whereas the remaining five neurons release neurotransmitters that produce IPSPs in the post-synaptic cells. When each of the twenty neurons is stimulated, it release enough neurotransmitters to produce a 2mV change (either positive or negative) in the postsynaptic membrane potential. If the threshold of the post-synaptic cell is 10mV greater than its resting potential, how many excitatory neurons must be fired if all five inhibitor neurons are stimulated (assume that spatial summation occurs). Also, what is the significance of this summation related to physiological changes.
Answer #1

Answer:

  • The should of postsynaptic neuron for excitation = 10mV above resting potential.
  • So the net effect of all the potentials breeds to increase the membrane potential by 10 milivolts. If all the Inhibitory.
  • Signals are activated, they cause further decrease in membrane potential by 2*5= 10 milivolts.
  • Hence the total positive signals by the EPSPs required is 20 milivolts, to neutralises the effect of the IPSPs and to depolarise the membrane.
  • Since each neuron generates 2mV, and there is spatial summation, number of neurons which need to be stimulated= 20/2 = 10.
  • So, minimum of 10 of the excitatory neurons need to be stimulated to cause action potential.

                                                                Please Rate My Answer.......Thank.......u...